OpenStudy (anonymous):

a^2-b^2/a^3+b^3 divided by 8b-8a/9 a+b/a^2+ab+b^2 as my answer but I don't think its right

OpenStudy (anonymous):

Show your work, please.

OpenStudy (paxpolaris):

\[{a^2-b^2 \over a^3+b^3} \div {8b-8a \over9}\] ?

OpenStudy (paxpolaris):

\[={a^2-b^2 \over a^3+b^3} \times {9\over 8b-8a}\] \[={(a+b)(a-b) \over (a+b)(a^2-ab+b^2)} \times {9\over -8(a-b)}\]

OpenStudy (paxpolaris):

\[\large{=-{9 \over 8(a^2-ab+b^2)} }\]